### v1

# Gaussian Generating functionals on easy quantum groups

Abstract.
We describe all Gaussian generating functionals on several easy quantum groups given by non-crossing partitions. This includes in particular the free unitary, orthogonal and symplectic quantum groups. We further characterize central Gaussian generating functionals and describe a centralization procedure yielding interesting (non-Gaussian) generating functionals.
Key words and phrases:
Gaussian functionals; partition quantum groups; quantum Lévy processes
2010 Mathematics Subject Classification:
Primary: 46L53; Secondary 81R50
A.S. was partially supported by the National Science Centre (NCN) grant no. 2020/39/I/ST1/01566 and by a grant from the Simons Foundation. A.S. would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Quantum Information, Quantum Groups and Operator Algebras, where work on this paper was completed. This was supported by EPSRC grant no EP/R014604/1.
1. Introduction
The study of stochastic processes on (locally) compact groups naturally focuses primarily on the most natural class, that of Lévy processes, i.e. stationary processes with independent, identically distributed increments. These can be naturally described via the associated convolution semigroups of probability measures, and further via their stochastic generators, which can be fully classified via the Lévy-Khintchin formula in its various incarnations ([Hey], [Lia]).
The arrival of quantum probability, and parallel developments related to quantum group theory, led in the 1980s to the emergence of the theory of quantum Lévy processes [ASW]. They can be studied from a purely algebraic point of view, as – similarly to their classical counterparts – all the stochastic information they carry is contained in the associated (quantum) convolution semigroup, which this time is a family of states on the underlying *-bialgebra, or in its generating functional ([Sch2]).
Among classical Lévy processes the most important are the Gaussian processes. The desire to understand their quantum equivalents led Schürmann to introduce the notion of a Gaussian generating functional. These can be in fact defined and studied on any augmented algebra (i.e. a complex unital *-algebra $A$ equipped with a character $\varepsilon:A\to\mathbb{C}$). We will however mostly focus on the original context of the algebras associated with compact quantum groups in the sense of Woronowicz, denoted below $\textup{Pol}(\mathbb{G})$, and further simply speak of Gaussian generating functionals on $\mathbb{G}$, where $\mathbb{G}$ is a given compact quantum group. The study of Gaussian generating functionals has many interesting connections to general quantum stochastic considerations ([DaG]), to Lévy-Khintchin decompositions and cohomology questions ([DFKS]) or to determining Gaussian parts of certain quantum groups ([FFS]). In this article we will focus on the classification problems for Gaussian functionals on certain free/universal quantum groups ([BaS]). The main result of our work can be summarised as the following statement.
Theorem.
Every Gaussian generating functional $\phi$ on the free unitary quantum group $U_{N}^{+}$ admits a unique decomposition into the sum of a “drift” part $D_{H}$ determined by an anti-hermitian matrix $H\in M_{N}(\mathbb{C})$ and a “diffusion” part $\Gamma_{W}$ determined by a matrix $W\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})$ satisfying a certain positivity condition. Conversely, every pair $H,W$ as above leads to a Gaussian generating functional $\phi=D_{H}+\Gamma_{W}$.
The result, which has obvious classical analogues, allows us further to obtain analogous descriptions for free orthogonal groups, for free symplectic group and for certain other classes of easy/partition quantum groups. One should note that the theorem above might be helpful in solving the outstanding problem left open in [FFS]: do Gaussian processes “see all of $U_{N}^{+}$”? Or, formally speaking, is $U_{N}^{+}$ its own Gaussian part?
A particularly interesting class of quantum convolution semigroups (or more generally quantum probability measures) on compact quantum groups is given by central states (see for example [CFK] or [FSW]). Motivated by this fact, we describe all central Gaussian generating functionals on the quantum groups listed above. These turn out to be rather limited; for example the free orthogonal group does not admit any central Gaussian processe. We show however how to produce central stochastic generators which are not Gaussian, but from a certain perspective can be viewed as quantum analogues of classical Brownian motions. These have been very recently studied in [Del], with natural cut-off estimates obtained.
The detailed plan of the paper is as follows: in Section 2 we recall the basic notions and background results we need to study Gaussian generating functionals. Section 3 treats the free unitary group and establishes the key theorem mentioned above. There we also discuss the Gaussian generating functionals on the infinite quantum hyperoctahedral group and on the duals of classical free groups. In Section 4 we consider free orthogonal group and its symplectic counterpart. Section 5 is devoted to the analysis of central Gaussian functionals. We first show that they appear relatively rarely, and then we describe the centralizing procedure and apply it to Gaussian generators.
2. Preliminaries
In this preliminary section, we recall some basic definitions and facts regarding Gaussian generating functionals.
The basic object of our study is an augmented unital complex *-algebra, i.e. a pair $(B,\varepsilon)$, where $B$ is a unital *-algebra and $\varepsilon:B\to\mathbb{C}$ a character (unital *-homomorphism). For simplicity we will always call such a pair an augmented algebra. A lot of the motivation for our study comes from the examples of augmented algebras given by the Hopf *-algebra $\textup{Pol}(\mathbb{G})$ of a compact quantum group $\mathbb{G}$ equipped with the counit, but we will stick for the moment to the general setting. A generating functional on $B$ is a linear functional $\phi:B\to\mathbb{C}$ with the following three properties:
  1. (1)
    $\phi(\mathbf{1})=0$ (normalization);
  2. (2)
    $\phi(b^{*})=\overline{\phi(b)}$ for all $b\in B$ (hermitianity);
  3. (3)
    $\phi(b^{*}b)\geqslant 0$ for all $b\in\ker(\varepsilon)$ (conditional positivity).
We are interested in generating functionals, because it follows from [Sch2, Section 3.2] that if $B$ happens to be a *-bialgebra, such functionals are in one-to-one correspondence with convolution semigroups of states (see for example [DFKS] for more information on the topic).
We now turn to the notion of Gaussianity for a generating functional. This is expressed through specific ideals of an augmented *-algebra $B$ which we now define. Set $K_{1}(B)=\ker(\varepsilon)$ and
$K_{n}(B)=\mathrm{Span}\{b_{1}\cdots b_{n}\mid b_{1},\cdots,b_{n}\in\ker(\varepsilon)\}=K_{1}(B)^{n}.$
Definition 2.1.
A generating functional $\phi:B\to\mathbb{C}$ on an augmented *-algebra $B$ is called Gaussian (or quadratic, [Sch2, Section 5.1]), if $\phi_{\mid K_{3}(B)}=0$.
Note that the defining property of Gaussian generating functionals translates into the following condition, valid for all $a,b,c\in B$:
$\phi(abc)=\phi(ab)\varepsilon(c)+\phi(ac)\varepsilon(b)+\phi(bc)\varepsilon(a)-\phi(a)\varepsilon(bc)-\phi(b)\varepsilon(ac)-\phi(c)\varepsilon(ab).$
This gives an inductive algorithm to compute $\phi$, see [FFS, Prop 2.7]. Note also that Gaussian generating functionals form a cone inside $B^{\prime}$. We need two more algebraic definitions.
Definition 2.2.
Let $(B,\varepsilon)$ be an augmented *-algebra and let $V$ be a vector space. A linear map $\eta:B\to V$ is called a Gaussian cocycle (or an $\varepsilon$-derivation) if for all $a,b\in B$,
$\eta(ab)=\varepsilon(a)\eta(b)+\eta(a)\varepsilon(b).$
Definition 2.3.
Given an augmented algebra $(B,\varepsilon)$ and a functional $\psi:B\to\mathbb{C}$ we define the coboundary of $\psi$, namely a functional $\partial\psi:B\otimes B\to\mathbb{C}$, by the (linear extension of the) formula
(2.1) $\partial\psi(a\otimes b)=\psi(ab)-\varepsilon(a)\psi(b)-\psi(a)\varepsilon(b),$
for $a,b\in B$.
A version of the GNS construction, starting from a (Gaussian) generating functional shows the following facts (contained in [Sch2, Subsection 2.3] and [Sch2, Proposition 5.1.1]).
Theorem 2.4.
Let $(B,\varepsilon)$ be an augmented *-algebra and let $\phi:B\to\mathbb{C}$ be a linear functional. The following facts are equivalent:
  • (i)
    $\phi$ is a Gaussian generating functional;
  • (ii)
    $\phi$ is hermitian and there exist a pre-Hilbert space $D$ and a Gaussian cocycle $\eta:B\to D$ such that for all $a,b\in B$,
    (2.2) $\partial\phi(a^{*}\otimes b)=\langle\eta(a),\eta(b)\rangle.$
We can moreover assume that $\eta(B)=D$ (the cocycle $\eta$ is then called surjective). Given two surjective cocycles $\eta:B\to D,\eta^{\prime}:B\to D^{\prime}$ such that (2.2) holds for both $\eta$ and $\eta^{\prime}$, we have a natural unitary equivalence between $\eta$ and $\eta^{\prime}$.
Definition 2.5.
A pair $(\phi,\eta)$, where $\phi:B\to\mathbb{C}$ is a Gaussian generating functional and $\eta:B\to D$ is a Gaussian cocycle such that (2.2) holds, is called a Gaussian pair. Given a Gaussian cocycle $\eta:B\to D$ we will say that it admits a Gaussian generating functional if there exists a Gaussian generating functional $\phi:B\to\mathbb{C}$ such that $(\phi,\eta)$ form a Gaussian pair.
Remark.
A special case of Gaussian pairs is that given by these of the form $(\phi,0)$. These arise from drifts $\phi$: hermitian linear functionals which are at the same time $\varepsilon$-derivations. Note that drifts are Gaussian functionals, and given a Gaussian pair $(\phi,\eta)$ and a functional $\psi\in B^{\prime}$ the pair $(\psi,\eta)$ is Gaussian if and only if $\psi-\phi$ is a drift.
The next proposition will be very useful in Section 4.
Proposition 2.6.
Assume that $(B,\varepsilon)$ and $(\tilde{B},\tilde{\varepsilon})$ are augmented *-algebras and that $q:\tilde{B}\to B$ is a morphism of augmented algebras, i.e. a unital *-homomorphism such that $\varepsilon\circ q=\tilde{\varepsilon}$. Then any Gaussian generating functional $\phi:B\to\mathbb{C}$ induces a Gaussian generating functional $\tilde{\phi}=\phi\circ q$ on $\tilde{B}$. If moreover $q$ is surjective, then there is a natural bijective correspondence between Gaussian pairs on $B$ and those Gaussian pairs on $\tilde{B}$ which vanish on $\ker(q)$.
Proof.
This follows directly from the fact that $q(K_{n}(\tilde{B}))\subset K_{n}(B)$ for all $n\in\mathbb{N}$. ∎
To check that a Gaussian generating functional vanishes on a certain ideal determined by a set of relations, we will rely on the following result proved in [FFS, Cor 2.8] in the language of Hopf *-algebras of compact quantum groups. Here we will state it in a formally broader context of augmented *-algebras, but the proof remains exactly the same.
Lemma 2.7.
Let $(B,\varepsilon)$ be an augmented *-algebra and let $\phi:B\to\mathbb{C}$ be a Gaussian generating functional. Assume that we have two families $\mathcal{X}=\{a_{1},\cdots,a_{n}\}\subset B$ and $\mathcal{Y}=\{b_{1},\cdots,b_{m}\}\subset\ker(\varepsilon)$ such that
  • (1)
    the family $\mathcal{X}$ generates $B$ as an algebra;
  • (2)
    $0=\phi(b_{k})=\phi(a_{j}b_{k})=\phi(b_{k}a_{j})$ for all $j=1,\cdots,n$, $k=1,\cdots,m$.
Then $\phi$ vanishes on the ideal generated by $\mathcal{Y}$.
Our goal in the sequel is to completely classify Gaussian functionals on augmented algebras associated with concrete compact quantum groups. The specific quantum groups that we will study are called free easy quantum groups and were introduced under that name in [BaS]. However, we will not need their general definition, because it will be more convenient to work with a specific description in each case. Moreover, there are many of these for which the problem has already been solved. Indeed, it was proven in [FFS, Prop 4.10] that if the Hopf *-algebra $\textup{Pol}(\mathbb{G})$ of a compact quantum group is generated by projections, then $\textup{Pol}(\mathbb{G})$ admits no non-zero Gaussian functionals. This is in particular the case for the quantum permutation groups $S_{N}^{+}$ (see [Wan2] for the definition) and the quantum reflection groups $H_{N}^{s+}$ for $1\leqslant s<\infty$ (see [BaV] for the definition). Note also that we will often simply speak of Gaussian pairs or Gaussian functionals on $\mathbb{G}$ (as opposed to on $\textup{Pol}(\mathbb{G})$), as mentioned already in the introduction.
3. Free unitary quantum groups
We start our study with free unitary quantum groups, which were originally defined in [Wan1]. We refer the reader to [Fre] for a detailed treatment of the theory and the definitions of the objects that we will use, but most of our computations only involve the defining relations of the corresponding *-algebras, which we now give. Throughout this section we fix $N\in\mathbb{N}$.
Definition 3.1.
Let $\textup{Pol}(U_{N}^{+})$ be the universal *-algebra generated by $N^{2}$ elements $(u_{ij})_{1\leqslant i,j\leqslant N}$ such that the matrices $U=(u_{ij})_{i,j=1}^{N}$ and $\overline{U}=(u_{ij}^{*})_{i,j=1}^{N}$ are both unitary. It is easy to check that the formula $\varepsilon(u_{ij})=\delta_{ij}$, for all $1\leqslant i,j\leqslant N$ determines a character on $\textup{Pol}(U_{N}^{+})$.
To classify Gaussian pairs on $U_{N}^{+}$, we will first record an upgraded version of an observation contained already in [DFKS].
Lemma 3.2.
Any matrix $A=(a_{ij})_{1\leqslant i,j\leqslant N}\in M_{N}(\mathbb{C})$ determines a (unique) $\varepsilon$-derivation $D_{A}:\textup{Pol}(U_{N}^{+})\to\mathbb{C}$ through the formula
$D_{A}(u_{ij})=a_{ij}\text{ for all }1\leqslant i,j=1\leqslant N.$
Moreover we have
$D_{A}(u_{ij}^{*})=-a_{ji}\text{ for all }1\leqslant i,j=1\leqslant N.$
Thus, $D_{A}$ is a drift if and only if $A=-A^{*}$.
Proof.
The first statement is a special case of [DFKS, Prop 3.2] (with $R=\mathrm{I}_{N})$. The second is a consequence of the formula $\eta\circ S=-\eta$, valid for any Gaussian cocycle on a Hopf *-algebra associated with a compact quantum group, with $S$ denoting the antipode (see the proof of [FFS, Thm 3.11]), and the fact that $S(u_{ij})=u_{ji}^{*}$ for all $1\leqslant i,j\leqslant N$. The last statement follows from the fact that the complex conjugate of a $\mathbb{C}$-valued $\varepsilon$-derivation is also an $\varepsilon$-derivation and the injectivity of the map $A\mapsto D_{A}$. ∎
The following is the first version of the main result of this section.
Theorem 3.3.
Let $d\in\mathbb{N}$, let $L_{1},\cdots,L_{d}\in M_{N}(\mathbb{C})$ be such that
(3.1) $\sum_{r=1}^{d}L^{*}_{r}L_{r}=\sum_{r=1}^{d}L_{r}L^{*}_{r}$
and let $H\in M_{N}(\mathbb{C})$ be anti-hermitian ($H=-H^{*}$). Denote by $(e_{1},\cdots,e_{d})$ the usual orthonormal basis of $\mathbb{C}^{d}$. Then there exists a Gaussian cocycle $\eta:\textup{Pol}(U_{N}^{+})\to\mathbb{C}^{d}$ and a hermitian functional $\Gamma:\textup{Pol}(U_{N}^{+})\to\mathbb{C}$ such that for every $1\leqslant i,j\leqslant N$ we have
$\eta(u_{ij})=\sum_{r=1}^{d}(L_{r})_{ij}e_{r}\quad\&\quad\Gamma(u_{ij})=-\frac{1}{2}\sum_{r=1}^{d}(L_{r}^{*}L_{r})_{ij},$
and for all $a,b\in\textup{Pol}(U_{N}^{+})$,
$\partial{\Gamma}(a^{*}\otimes b)=\langle\eta(a),\eta(b)\rangle.$
The conditions above determine the pair $(\Gamma,\eta)$ uniquely. Set $\phi=\Gamma+D_{H}$. Then $(\phi,\eta)$ is a Gaussian pair, and moreover all Gaussian pairs with surjective cocycles on $\textup{Pol}(U_{N}^{+})$ (hence also all Gaussian generating functionals) arise in this way. We may in addition choose the matrices $L_{1},\cdots,L_{d}$ to be linearly independent.
Proof.
Assume first that we are given matrices $L_{1},\cdots,L_{d}\in M_{N}(\mathbb{C})$ satisfying Equation (3.1) and $H\in M_{N}(\mathbb{C})$ such that $H=-H^{*}$. Lemma 3.2 guarantees that for every $1\leqslant r\leqslant d$, we have an $\varepsilon$-derivation $D_{L_{r}}:\textup{Pol}(U_{N}^{+})\to\mathbb{C}$; it is then immediate that
$\eta:=\sum_{r=1}^{d}D_{L_{r}}e_{r}:\textup{Pol}(U_{N}^{+})\to\mathbb{C}^{d}$
is also an $\varepsilon$-derivation.
Let us introduce matrices $B,\tilde{B}\in M_{N}(\mathbb{C})$ through the formulæ (for $1\leqslant i,j\leqslant N$)
(3.2) $B_{ij}=\sum_{p=1}^{N}\langle\eta(u_{ip}),\eta(u_{jp})\rangle=\sum_{r=1}^{d}\sum_{p=1}^{N}\overline{(L_{r})_{ip}}(L_{r})_{jp}=\sum_{r=1}^{d}(L_{r}L_{r}^{*})_{ji},$
(3.3) $\tilde{B}_{ij}=\sum_{p=1}^{N}\langle\eta(u_{ip}^{*}),\eta(u_{jp}^{*})\rangle=\sum_{r=1}^{d}\sum_{p=1}^{N}\overline{(L_{r})_{pi}}(L_{r})_{pj}=\sum_{r=1}^{d}(L_{r}^{*}L_{r})_{ij},$
where in the second string of equalities we use Lemma 3.2.
The forward direction of [DFKS, Theorem 3.3] states that if $B=\tilde{B}^{t}$ (which as the above computation shows is equivalent to Equation (3.1)), then $\eta$ admits a generating functional. Moreover the generating functional $\Gamma$ constructed in the proof of [DFKS, Theorem 3.3] satisfies for all $1\leqslant i,j\leqslant N$ the equalities
(3.4) $\Gamma(u_{ij})=-\frac{1}{2}\tilde{B}_{ij}=-\frac{1}{2}\sum_{r=1}^{d}(L_{r}^{*}L_{r})_{ij},$
which is the condition displayed in the statement. The uniqueness claim follows as $\Gamma$ is assumed to be hermitian, so that the algebraic conditions determine both $\eta$ and $\Gamma$ in terms of their values on generators $u_{ij}$. Eventually, the last part of Lemma 3.2 shows that $D_{H}$ is a drift. Thus $(\phi,\eta)$ is a Gaussian pair.
Assume conversely that $\phi$ is a Gaussian functional on $\textup{Pol}(U_{N}^{+})$. By Theorem 2.4, we can assume that it is a part of a Gaussian pair $(\phi,\eta)$, where $\eta:\textup{Pol}(U_{N}^{+})\to D$ is surjective. As the derivation property implies that the image of $\eta$ is spanned by the images of the generators $u_{ij}$ for $1\leqslant i,j\leqslant N$, the space $D$ must be finite dimensional. Set $d=\dim(D)$ and identify $D$ with $\mathbb{C}^{d}$. Then, set for $1\leqslant r\leqslant d$, $1\leqslant i,j\leqslant N$
$(L_{r})_{i,j}=\langle e_{r},\eta(u_{ij})\rangle,$
so that
$\eta=\sum_{r=1}^{d}D_{L_{r}}e_{r}.$
The argument in the first part of the proof shows that $\eta$ admits a generating functional $\Gamma:\textup{Pol}(U_{N}^{+})\to\mathbb{C}$ satisfying the conditions listed in the theorem. As $(\Gamma,\eta)$ is a Gaussian pair, by the remarks after Definition 2.5 we must have $\Gamma=\phi+\omega$, where $\omega:\textup{Pol}(U_{N}^{+})\to\mathbb{C}$ is a drift. Thus the last part of Lemma 3.2 ends the proof of the main statement of the theorem.
Eventually, let us prove that the cocycle
$\eta=\sum_{r=1}^{d}D_{L_{r}}$
is surjective if and only if the matrices $L_{1},\cdots,L_{d}$ are linearly independent. Indeed, as mentioned above, the range of $\eta$ is spanned by the elements $\eta(u_{ij})$, and it is easy to check that given a vector $\xi=(\lambda_{1},\cdots,\lambda_{d})\in\mathbb{C}^{d}$, we have
$\left(\langle\xi,\eta(u_{ij}\rangle\right)_{i,j=1}^{N}=\sum_{r=1}^{d}\overline{\lambda_{r}}L_{r},$
hence $\xi\perp\mathrm{Ran}(\eta)$ if and only if $\sum\overline{\lambda_{r}}L_{r}=0$, and the result follows. ∎
Note that given a fixed anti-hermitian $H\in M_{N}(\mathbb{C})$, several tuples of linearly independent matrices $(L_{1},\cdots,L_{d})$ as above can yield the same generating functional (and equivalent Gaussian pairs). The proof above shows that the freedom is related to the choice of the orthonormal basis in the carrier space of the cocycle $\eta$, which affects the matrices $L_{r}$ via a unitary transformation. In other words, given an alternative tuple of linearly independent matrices $(\tilde{L}_{1},\cdots,\tilde{L}_{d})$ leading – together with $H$ – to the same $\phi$, we have a unitary $U\in M_{d}(\mathbb{C})$ such that for all $1\leqslant i\leqslant d$, $\tilde{L}_{i}=\sum_{r=1}^{d}U_{ri}L_{r}$. Note that this operation does not affect the matrix
$W=\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*}\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C}).$
This observation leads to the second main theorem of this section. We begin with a lemma characterising the matrices of the form mentioned above.
Lemma 3.4.
Consider a matrix $W\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})$,
$W=\sum_{i,j,k,l=1}^{n}W_{ij,kl}e_{ij}\otimes e_{kl}$
for certain coefficients $W_{ij,kl}\in\mathbb{C}$. Then, the following conditions are equivalent:
  • (i)
    there exists $d\in\mathbb{N}$ and matrices $L_{1},\cdots,L_{d}\in M_{N}(\mathbb{C})$ such that
    $W=\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*};$
  • (ii)
    for every matrix $X\in M_{N}(\mathbb{C})$, we have
    $\sum_{i,j,k,l=1}^{N}\overline{X_{ik}}W_{ki,jl}X_{jl}\geqslant 0.$
Proof.
Let $A,B\in M_{N}(\mathbb{C})$ and define a map $\Psi_{A\otimes B}:M_{N}(\mathbb{C})\to M_{N}(\mathbb{C})$ by the formula
$\Psi_{A\otimes B}(Z)=AZB\text{ for }Z\in M_{N}(\mathbb{C}).$
This extends by bilinearity to a linear map $\Psi_{W}:M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})\to B(M_{N}(\mathbb{C}))$ which is easily seen to be injective. Applying the Kraus characterisation of completely positive maps on matrices, we see that $\Psi_{W}$ is completely positive if and only if $W$ satisfies (i). On the other hand, by Choi’s theorem, $\Psi_{W}$ is completely positive if and only if its $N$-th matrix lifting maps the Choi matrix $E:=(e_{ij})_{i,j=1}^{N}\in M_{N}(M_{N}(\mathbb{C}))$ to a positive matrix. We therefore compute:
$\Psi_{W}^{(N)}(E)=(\Psi_{W}(e_{ij}))_{i,j=1}^{n}=\left(\sum_{k,l=1}^{N}W_{ki,jl}e_{kl}\right)_{i,j=1}^{N},$
so that $\Psi_{W}^{(N)}(E)\geqslant 0$ if and only if for any vectors $\xi_{1},\cdots,\xi_{N}\in\mathbb{C}^{N}$ we have
$\sum_{i,j=1}^{N}\left\langle\xi_{i},\sum_{k,l=1}^{N}W_{ki,jl}e_{kl}\xi_{j}\right\rangle\geqslant 0,$
which is precisely condition (ii) in the lemma (set $X_{ij}=\langle e_{i},\xi_{j}\rangle$ for $1\leqslant i,j\leqslant N$). ∎
Remark.
Condition (3.1) can be expressed in terms of
$W=\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*}$
simply as $M(W)=M(\Sigma(W))$, where $M$ denotes the multiplication of two elements of the tensor product, and $\Sigma$ is the tensor flip. We can also express it using the map $\Psi_{W}$. Indeed, viewing $M_{N}(\mathbb{C})$ as a Hilbert space, when equipped with the scalar product $\langle X,Y\rangle:=\mathrm{Tr}(X^{*}Y)$ (for $X,Y\in M_{N}(\mathbb{C})$), given a map $\Psi:M_{N}(\mathbb{C})\to M_{N}(\mathbb{C})$ we can also consider its Hilbert space adjoint $\Psi^{*}:M_{N}(\mathbb{C})\to M_{N}(\mathbb{C})$. Assume then that $W\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})$ decomposes as $W=\sum_{i=1}^{r}L_{i}\otimes L_{i}^{*}$. Then, for all $X,Y\in M_{N}(\mathbb{C})$,
$\langle X,\Psi_{W}(Y)\rangle=\mathrm{Tr}\left(X^{*}\sum_{i=1}^{d}L_{i}YL_{i}^{*}\right)=\sum_{i=1}^{r}\mathrm{Tr}(L_{i}^{*}X^{*}L_{i}Y)$
so that $(\Psi_{W})^{*}=\Psi_{\Sigma(W)}$. Thus, $M(W)=M(\Sigma(W))$ if and only if $\Psi_{W}(1)=\Psi_{W}^{*}(1)$.
We can now give the second form of our classification of Gaussian generating functionals on free unitary quantum groups.
Theorem 3.5.
There is a one-to-one correspondence between Gaussian generating functionals on $\textup{Pol}(U_{N}^{+})$ and pairs $(W,H)$ where
  1. (1)
    $H\in M_{N}(\mathbb{C})$ is anti-hermitian;
  2. (2)
    $W=(w_{ijkl})_{1\leqslant i,j,k,l\leqslant N}\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})$ satisfies $M(W)=M(\Sigma(W))$ and the positivity condition:
    $\sum_{i,j,k,l=1}^{N}\overline{X_{ik}}W_{ki,jl}X_{jl}\geqslant 0.$
    for all $X=(X_{jk})_{1\leqslant j,k\leqslant N}\in M_{N}(\mathbb{C})$.
Proof.
Given the pair $(W,H)$ as above we can use Lemma 3.4 to obtain matrices $L_{1},\cdots,L_{d}\in M_{N}(\mathbb{C})$ such that
$W=\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*};$
as $M(W)=M(\Sigma(W))$, Condition (3.1) holds and we can apply Theorem 3.3 to obtain a Gaussian generating functional $\phi=\Gamma+D_{H}$.
Conversely, given a Gaussian generating functional $\phi$, we obtain by Theorem 3.3 an anti-hermitian matrix $H$ and matrices $L_{1},\cdots,L_{d}\in M_{N}(\mathbb{C})$ such that $\phi=\Gamma+D_{H}$. We can simply set $W=\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*}$.
It remains to show that the correspondence described above is bijective; in other words, that given a Gaussian generating functional $\phi$ we can determine $(W,H)$. This could be deduced from the proof of Theorem 3.3 and remarks after the theorem, but we can also argue directly. Assume that $\phi=\Gamma+D_{H}$ is constructed as in the first part of the proof. Then, $\Gamma$ is hermitian and by the formulæ (3.2)-(3.4), we have for each $1\leqslant i,j\leqslant N$
(3.5) $\Gamma(u_{ij})=-\frac{1}{2}\sum_{r=1}^{d}\sum_{p=1}^{N}\overline{(L_{r})_{pi}}(L_{r})_{pj}=\overline{-\frac{1}{2}\sum_{r=1}^{d}\sum_{p=1}^{N}\overline{(L_{r})_{pj}}(L_{r})_{pi}}=\overline{\Gamma(u_{ji})}=\Gamma(u_{ji}^{*}).$
On the other hand
$D_{H}(u_{ij})-D_{H}(u_{ji}^{*})=H_{ij}-\overline{D_{H}(u_{ji})}=H_{ij}-\overline{H_{ji}}=H_{ij}-(H^{*})_{ij}=2H_{ij}.$
Thus for every $1\leqslant i,j\leqslant N$ we have
(3.6) $H_{ij}=\frac{1}{2}(\phi(u_{ij})-\phi(u_{ji}^{*})).$
Let then $\eta:\textup{Pol}(U_{N}^{+})\to\mathbb{C}^{d}$ be a surjective cocycle such that $(\phi,\eta)$ is a Gaussian pair, and recall that we can write
$\eta=\sum_{r=1}^{d}L_{r}.$
Fix $i,j,k,l\in\{1,\cdots,N\}$ and consider the following computation:
$\partial\phi(u_{ij}\otimes u_{kl})$ $=\langle\eta(u_{ij}^{*}),\eta(u_{kl})\rangle=-\langle\eta(u_{ji}),\eta(u_{kl})\rangle=-\sum_{r=1}^{d}\overline{(L_{r})_{ji}}(L_{r})_{kl}$
$=-\left(\sum_{r=1}^{d}L_{r}\otimes L_{r}^{*}\right)_{(k,i),(l,j)},$
so that in the end,
(3.7) $\partial\phi(u_{ij}\otimes u_{kl})=-W_{ki,lj}.$
Definition 3.6.
Given matrices $W\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})$ and $H\in M_{N}(\mathbb{C})$ satisfying the conditions stated in Theorem 3.5 we will denote the Gaussian generating functional associated to $W$ and 00 by $\Gamma_{W}$, so that any Gaussian generating functional $\phi$ on $\textup{Pol}(U_{N}^{+})$ decomposes uniquely as
$\phi=\Gamma_{W}+D_{H}.$
We will then call $D_{H}$ the drift part of $\phi$ and $\Gamma_{W}$ the diffusion part of $\phi$.
Remark.
The space of anti-hermitian $N$ by $N$ complex matrices is nothing but the Lie algebra $\mathfrak{u}_{N}$ of the classical unitary group $U_{N}$. It is therefore in one-to-one correspondance with drifts on $U_{N}^{+}$. Moreover, one easily checks that if $\ast$ denotes the convolution of linear forms (induced by the coproduct), then
$D_{H}\ast D_{K}-D_{K}\ast D_{H}=D_{[H,K]}$
so that drifts form a Lie algebra isomorphic to $\mathfrak{u}_{N}$.
Remark.
Note that the formulæ (3.6)-(3.7) can be expressed via matrix liftings of $\phi$: if we write $U=(u_{ij})_{1\leqslant i,j\leqslant N}\in M_{N}(\textup{Pol}(U_{N}^{+}))$, and define $U\tilde{\otimes}U\in M_{N}(\mathbb{C})\otimes M_{N}(\mathbb{C})\otimes\textup{Pol}(U_{N}^{+})$ to be the matrix with entries $(U\tilde{\otimes}U)_{(i,k),(j,l)}=u_{ij}u_{kl}$, then we have
$H$ $=\frac{1}{2}\left(\phi^{(N)}(U)-\phi^{(N)}(U^{*}\right);$
$-W^{t\otimes t}$ $=\phi^{(N^{2})}(U\tilde{\otimes}U)-\phi^{(N)}(U)\otimes I_{N}-I_{N}\otimes\phi^{(N)}(U);$
$\phi^{(N)}(U)$ $=-\frac{1}{2}M(W)+H.$
We will now briefly discuss a possible characterisation of the decomposition $\phi=\Gamma_{W}+D_{H}$. We know which generating functionals are drifts, but there is no general notion of a ‘driftless’ (Gaussian) generating functional. But we have the following definition, writing for simplicity $K_{i}:=K_{i}(\textup{Pol}(U_{N}^{+}))$ with $i=1,2$.
Definition 3.7.
[…]